OpenStudy (anonymous):

Simplify. Assume that no denominator is 0 and that 0^0 is not considered. (4x/3) ^3

OpenStudy (anonymous):

(4x/3)*(4x/3)*(4x/3) 4x * 4x * 4x 4*4*4*x*x*x ---------- = ------------ 3 * 3 * 3 3*3*3 can you do the rest?

OpenStudy (anonymous):

Yes. I had (4x/3) ^3 4x *4x * 4x= 64x 3 * 3 * 3 = 27 64x/ 27 but I am unsure if this looks correct?

OpenStudy (anonymous):

\[\left(\frac{4 x}{3}\right)^3\text{=}\frac{64 x^3}{27} \]

OpenStudy (anonymous):

4*4*4 = 64 x*x*x = ??? 3*3*3 = ???

OpenStudy (anonymous):

Thank you!!